### v1

# Implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model in the GENIE neutrino event generator

Abstract
We discuss the first implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model for quasielastic (1p1h) and multinucleon (2p2h and 3p3h) excitations in the widely used GENIE neutrino event generator. Validation steps are presented, in particular, through direct comparisons of GENIE cross section output with original calculations performed by the authors of the model. Predictions for 12C, 16O and 40Ar are compared with some available T2K and MicroBooNE experimental measurements showing a reasonable agreement.
I Introduction
The success of the next generation long-baseline accelerator-based neutrino oscillation experiments Hyper-Kamiokande Abe et al. (2018a) and DUNE Acciarri et al. (2015) will depend on the reduction to the percent level of the systematic errors. Presently, the most important source of uncertainties is represented by the neutrino-nucleus cross sections Katori and Martini (2018); Alvarez-Ruso et al. (2018), due to the difficulty in modeling nuclear dynamics, the nucleus being a quantum many-body system whose constituents interact through the strong force. To improve the knowledge of neutrino cross sections, a common and joint effort of experimental measurements, theoretical calculations and model implementations in Monte Carlo event generators is needed. These three aspects cannot be separated from each other and are entangled as three Borromean rings. Think, for example, of the neutrino energy reconstruction problem, the background subtraction for a peculiar experimental signal topology, or the modeling of the detector response.
In this paper, we present the implementation of the theoretical RPA-based model of the Lyon group, described in detail in the article of Martini, Ericson, Chanfray and Marteau Martini et al. (2009) into the GENIE event generator Andreopoulos et al. (2010). We then compare its predictions with different experimental measurements. This model allows a unified description of several scattering channels: quasielastic, multinucleon emission and the coherent and incoherent single pion production. Within this model, the debated MiniBooNE quasielastic-like cross section on carbon Aguilar-Arevalo et al. (2010) has been explained for the first time Martini et al. (2009, 2010) by stressing the crucial role of the multinucleon (also called 2p2h or npnh) excitations contribution which sizably increase the cross section. After this explanation, 2p2h excitations started to attract a lot of attention (for a review see, for example, Refs. Katori and Martini (2018); Alvarez-Ruso et al. (2018)), first in the neutrino scattering community, then in the neutrino oscillation one. This growing interest is due to the central role of npnh, together with quasielastic processes, in the accurate reconstruction of the neutrino energy using the kinematic method employed by the T2K Abe et al. (2011) and Hyper-K experiments Martini et al. (2012, 2013); Nieves et al. (2012); Lalakulich et al. (2012); Ankowski et al. (2015), and consequently in the precise determination of the neutrino oscillation parameters. All widely-used event generators now include the multinucleon excitation channel, previously ignored before the MiniBooNE quasielastic-like results, into their packages. For a relatively recent state-of-the-art implementation in the different generators, see for example Refs. Alvarez-Ruso et al. (2021); Hayato and Pickering (2021); Golan et al. (2012); Buss et al. (2012), as well as the dedicated articles Schwehr et al. (2016); Dolan et al. (2020); Bourguille et al. (2021); Prasad et al. (2024). This implementation effort is still ongoing, essentially for two reasons. Firstly, various theoretical models present some differences in their predictions for this channel, whose complexity – stemming from the intrinsic difficulty of nuclear physics – makes it hard to treat without approximations. Secondly, the experimental neutrino community is moving towards more exclusive measurements based on detected hadronic variables, whereas most of the existing models were originally developed for inclusive cross sections in terms of leptonic observables.
Up until now, the model of Ref. Martini et al. (2009) has not been included in any generators by the implementation of the corresponding theoretical hadron tensor, but only employing tables of shared cross sections at different kinematics via a re-weighting procedure Nguyen (2022). The hadron tensor strategy is employed here for the first time for this model. For a detailed explanation of this strategy, see Ref. Dolan et al. (2020), where it is introduced in the context of the SuSAv2 model. Even though, in principle, all the channels described by this model can be implemented in the generators, in the present work we confine ourselves to considering only the quasielastic (also called 1p1h) and the multinucleon excitations channel (2p2h and 3p3h, collectively referred to as npnh), which have been the most investigated and employed in the neutrino analyses of accelerator-based experiments over the last fifteen years. For this purpose we follow the same protocol as the one adopted in Refs. Dolan et al. (2020) and Dolan et al. (2022) for the SuSAv2+MEC Megias et al. (2016) (quasielastic + 2p2h) and Ghent CRPA Pandey et al. (2015) (giant resonances + quasielastic) models, respectively.
The interest in implementing the Martini, Ericson, Chanfray, and Marteau model (hereafter referred to as the Martini et al. model) relies on several aspects. First, as already mentioned, differences appear among the various models, and none is currently able to simultaneously describe all the available experimental measurements across the full neutrino energy range – from low (e.g., T2K) to high (e.g., MINERvA) energies – and across nuclear targets (most notably, 12C, 16O and 40Ar). In the absence of such a solution, the span of the predictions from plausible models already implemented inside generators provides some guidance about, for example, the size of systematic errors associated to different processes (e.g., see the heavy use of this approach in Ref. Wilkinson et al. (2022)). Second, in the Martini et al. model the different 2p2h contributions – namely the ones related to i) nucleon-nucleon correlations, ii) meson exchange currents (MEC), and iii) their interference – as well as the 3p3h excitations, are separately calculated and available as distinct components of the hadron tensor. This allows for easier comparison between 2p2h models and, overall, could help avoid the risk of double counting (e.g., nucleon-nucleon correlations contributions) when combining different channels taken from different models (e.g., using a spectral function model, like that in Ref. Benhar and Meloni (2007), for the quasielastic channel and Fermi gas-based models for 2p2h interactions). Finally, calculations for different atomic nuclei are available, enabling the modeling of cross sections for carbon, water and argon targets – the main targets of present and future accelerator-based neutrino experiments.
The article is organized as follows. In Sec. II, we briefly recall the main aspects and achievements of the model presented in Ref. Martini et al. (2009). In Sec. III, we describe the strategy for implementing this model in GENIE and its validation. In Sec. IV, we compare the model predictions with several sets of measurements, with the aim of illustrating the model’s potential for describing different reaction channels, various final-state topologies, and different targets. Finally, in Sec. V, we draw our conclusions.
II The theoretical model
The theoretical model is described in Ref. Martini et al. (2009), to which we refer for all the details. It is based on nuclear response functions treated in the random phase approximation (RPA) on top of a local Fermi gas (LFG) calculation. This formalism belongs to a long tradition of investigations: first on electron scattering and pionic modes in nuclei (see, for example, Refs. Delorme et al. (1980); Alberico et al. (1980, 1982, 1984); Stroth et al. (1985); Alberico et al. (1985, 1986); Chanfray and Schuck (1988); Delorme and Guichon (1991); Laktineh et al. (1993); Chanfray et al. (1993); Alberico et al. (1998)), and later on neutrino scattering Delorme and Ericson (1985); Marteau (1999); Marteau et al. (2000). The model of Ref. Martini et al. (2009) is the continuation and extension of these works. One of its strengths is that it allows a unified description of several scattering channels: quasielastic, multinucleon emission and the coherent and incoherent single pion production. As already mentioned, within this model, the debated MiniBooNE quasielastic-like cross section on carbon Aguilar-Arevalo et al. (2010) was explained for the first time Martini et al. (2009, 2010) by advocating the crucial role of the multinucleon excitations. In Ref. Martini et al. (2011), an improvement of the original description of the quasielastic channel was made through the introduction of relativistic corrections. This relativistic description of the quasielastic channel is implemented in the present work. The Martini et al. model has successfully reproduced many of the features of several muon-neutrino flux-integrated double differential cross-section measurements in the past: the MiniBooNE Charged Current Quasielastic (CCQE) - like neutrino Martini et al. (2011) and antineutrino Martini and Ericson (2013) measurements, as well as their combination Ericson and Martini (2015); the MiniBooNE CC1 and the T2K CC inclusive measurements Martini and Ericson (2014); Martini et al. (2022); the T2K CC0 for neutrino Abe et al. (2016) and antineutrino measurements, as well as their combination Abe et al. (2020a). Single differential electron-neutrino T2K CC inclusive cross sections have been successfully reproduced as well Martini et al. (2016). More recently, the model has been employed to investigate the MicroBooNE CC inclusive Martini et al. (2022) and NC1 Martini et al. (2024a) cross sections on argon. The counter-intuitive dominance of the differential cross sections over the ones in specific kinematical conditions – not only for quasielastic, but also for multinucleon excitations and single pion production – has been explained with simple arguments in Ref. Martini et al. (2024b) using this model.
Even though it has not yet been included in any Monte Carlo event generator via the implementation of the corresponding hadron tensor, the model has already been considered in many experimental neutrino cross-section measurement comparisons Abe et al. (2016, 2020a); Acero et al. (2020); Abe et al. (2023a) and oscillation analyses Abe et al. (2017, 2018b, 2020b, 2023b, 2025). In these experimental papers, the model is often referred to as “Martini et al.” or “Lyon”. The main reason for considering it is its larger cross-section predictions in the 2p2h sector (especially for neutrinos, and to a lesser extent for antineutrinos) compared to those of Ref. Nieves et al. (2011), a similar theoretical approach, often referred to as “Nieves et al.” or “Valencia”, which is already implemented in many generators. A very recent reexamination Sobczyk and Nieves (2025) of the Nieves et al. approach now leads to new results that are closer to that of Martini et al.
Since the present work focuses mainly on multinucleon excitations, we recall that in the model of Ref. Martini et al. (2009) the following contributions are included: the nucleon-nucleon correlations (denoted NN in Ref. Martini et al. (2009)); the -MEC contribution (denoted in Ref.Martini et al. (2009)); and the NN correlations–-MEC interference, (denoted N in Ref. Martini et al. (2009)). We also recall that, as in all the works of the Lyon group starting from the end of Ref. Martini et al. (2009), only the npnh evaluation called “new” in Ref. Martini et al. (2009) is considered; the “old” version corresponds to the one employed in Refs. Marteau (1999); Marteau et al. (2000). This so-called “new” evaluation is based, for the NN and N parts, on the microscopic calculations of Ref. Alberico et al. (1984), aimed at determining the 2p2h contribution to the isospin spin-transverse response. For the part, the contributions related to the non-pionic decay of the inside the nuclear medium – leading to 2p2h and 3p3h excitations – are taken from Ref. Oset and Salcedo (1987), while the terms not reducible to a modifications of the width are taken from Ref. Delorme and Guichon (1991), which exploits the calculations of Ref. Shimizu and Faessler (1980). For further details, see Sec. IV C of Ref. Martini et al. (2009) and Sec. II of Ref. Martini et al. (2010). We also recall that all the form factors entering the hadron tensor coincide with those of Ref. Martini et al. (2009), with exactly the same parameters. For example, for the axial form factor, a standard dipole parameterization is adopted, with and .
III Implementation in GENIE and validation
The model described above has been implemented in the GENIE neutrino event generator using the same approach as that adopted in Refs. Dolan et al. (2020) and Dolan et al. (2022). This approach relies on the fact that the differential cross section can be expressed as the product of kinematic factors and the contraction of a generic lepton tensor with a model-specific hadron tensor. The latter encodes all of the nuclear dynamics. In this framework, the implementation of the Martini et al. model is achieved by inserting new hadron tensor look-up tables into GENIE, as was previously done for other models in Refs. Schwehr et al. (2016), Dolan et al. (2020) and Dolan et al. (2022). These tables are provided for the 1p1h and npnh channels for carbon, oxygen and calcium targets. The npnh contribution includes both 2p2h and 3p3h components, and separate hadron tensors are employed for each. Despite the naming, the 3p3h implementation does not produce three final-state nucleons; instead, it is treated as an additional 2p2h-like contribution and results in two nucleons in the final state. Limitations of this approach are discussed in Sec. III.2. The input hadron tensors are defined over an energy transfer varying in the range , and a momentum transfer varying in the range , as in Ref. Martini et al. (2009). They are binned ( bins in energy transfer and bins in momentum transfer) and are evaluated using the same interpolation method of Refs. Dolan et al. (2020) and Dolan et al. (2022), similar to the one described in Ref. Schwehr et al. (2016).
III.1 Validation and comparison of models
For validation purposes, the model predictions obtained via GENIE have been compared with the original calculations performed by the authors of the model Martini et al. (2009), including the relativistic corrections for the quasielastic, published in Ref. Martini et al. (2011) for neutrinos and in Ref. Martini and Ericson (2013) for antineutrinos. The corresponding charged-current cross sections on 12C as a function of (anti)neutrino energy are presented for and in Figs. 2 and 2, respectively (and in Figs. 10 and 10 of Appendix References for 16O and 40Ca). These plots demonstrate perfect agreement between the original calculations (shaded bands) and the GENIE predictions for the different contributions to the cross section: “1p1h” (dashed), “npnh” (dotted) and their sum (solid). The predictions of other 1p1h and 2p2h models available in GENIE (“Nieves et al.” and “SuSAv2”) are also shown for reference. The corresponding GENIE tunes used are G18_10a_00_000 for the Nieves et al. model and G21_11a_00_000 for SuSAv2. As already illustrated in other articles (see for example Fig. 9 of Ref. Abe et al. (2023a)) the npnh (as well as the 1p1h) contributions estimated by these three models are different, and their relative importance with respect to 1p1h varies between neutrinos and antineutrinos, as discussed in Refs. Martini et al. (2010); Katori and Martini (2018). It is also important to note that a one-to-one correspondence between different exclusive channel contributions can be misleading. For example the nucleon-nucleon short-range correlations are part of npnh channel in the Martini et al. model, whereas they are included in the 1p1h component for SuSAv2. As can be seen in Figs. 2 and 2, the sum of these two channels is similar for Martini et al. and SuSAv2, especially in the antineutrino case.
For the sake of comparison with other models, we show in Figs. 4 and 4 the T2K flux-integrated double differential cross section as a function of and for the 1p1h and npnh channels, corresponding to Fig. 4 of Dolan et al. (2022) and Fig. 2 of Dolan et al. (2020) respectively. In Fig. 4, one can easily recognize the quasielastic line
where denotes the nucleon mass. The spread around this curve is due to Fermi motion111The fact that the quasielastic region is not strictly limited by the two quasielastic bands , typical of a global relativistic Fermi gas-based model with Fermi momentum , is due to the local density approximation and to the presence of Bessel functions in the nuclear response definitions (see Appendix 2 of Ref. Martini et al. (2009)), which result in very small fluctuations beyond the quasielastic region.. From Fig. 4 one can observe that the multinucleon response region covers the whole available phase space from the input tensors. Furthermore, two distinct branches can be identified: one below the quasielastic region – absent in the SuSAv2 case, see Fig. 2 of Ref. Dolan et al. (2020) – arising from nucleon-nucleon correlations, and a second one above the quasielastic region, extending from the dip to the region, due to -MEC contributions222These two branches are more distinguishable looking directly to the phase space of the nuclear response function, rather than to the flux-integrated cross section. See for example Fig. 7 of Ref. Katori and Martini (2018).. Overall, these two branches are less clearly separated than in the Nieves et al. model (see again Fig. 2 of Ref. Dolan et al. (2020)); in the Martini et al. case, the cross section is concentrated between the two contributions, which add coherently together with their significant interference. To summarize, the phase-space behavior of the multinucleon cross section in the Martini et al. model differs from those of Nieves et al. and SuSAv2, though it shares features with both: the presence of two branches, as in Nieves et al., and the accumulation of cross section strength in the dip region, as in SuSAv2.
As a validation of the implementation, and to facilitate comparisons with other models, in Fig. 5 we present the double differential cross sections 333Here, denotes the solid angle element. Assuming azimuthal symmetry, the double differential cross section is related by . for both neutrino and antineutrino interactions on oxygen, at fixed incident energy and lepton scattering angle444GENIE does not allow the user to explicitly fix the scattering angle. Events with were selected to approximate a scattering angle of ., as a function of the energy transfer . This represents the most stringent validation test, as the cross sections are not integrated over any variable, in contrast to the total or flux-integrated differential cross sections discussed earlier. A perfect agreement is observed between the theoretical calculations (solid lines) and the GENIE implementation (histograms) for both 1p1h and npnh channels, and for neutrino as well as antineutrino interactions555In the 1p1h case, the small fluctuations at low discussed in footnote 1 can also be seen..
Fig. 5 can be directly compared with Fig. 9 of Ref. Sobczyk and Nieves (2025), the recent reexamination by Sobczyk and Nieves of the multinucleon cross-section model originally presented in Ref. Nieves et al. (2011). In Ref. Sobczyk and Nieves (2025), the revised npnh single differential cross section at fixed neutrino energy, as well as the MiniBooNE flux-integrated double differential cross section on carbon, are compared with the corresponding results of Martini et al. Martini et al. (2009, 2011). The updated Valencia predictions are shown to be closer to the Lyon model, thus ending a long-standing discrepancy between the two.
We now extend the comparison to the double differential cross section at fixed neutrino energy. Before focusing on the npnh channel, we note that the 1p1h distribution in Fig. 5 exhibits a shape similar to that in Fig. 9 of Ref. Sobczyk and Nieves (2025), as expected, since both are based on RPA calculations on top of a local relativistic Fermi gas. However, the distribution in Fig. 5 is slightly more peaked and shifted to lower by approximately . This is probably due to the absence of removal energy in the calculations of Martini et al. unlike the case of Nieves et al. These small differences are often washed out in flux-integrated double differential distributions (see, e.g., Fig. 10 of Ref. Sobczyk and Nieves (2025)).
Turning to the npnh channel, for neutrinos, the evaluation of Ref. Sobczyk and Nieves (2025) aligns more closely with the results of Ref. Martini et al. (2009) than with the earlier ones of Ref. Nieves et al. (2011). However, the double-bump structure shown in Fig. 9 of Ref. Sobczyk and Nieves (2025), attributed to nucleon-nucleon correlations and MEC contributions, is not visible in Fig. 5. For antineutrinos, the npnh contribution in Fig. 5 is smaller than that in Fig. 9 of Ref. Sobczyk and Nieves (2025); nonetheless, the total cross section as a function of antineutrino energy in Fig. 8 of Ref. Sobczyk and Nieves (2025) appears to be closer to the results of Ref. Martini et al. (2009) than to the previous Valencia model predictions shown in Fig. 2.
To summarize, although the revised npnh model of Valencia reported in Ref. Sobczyk and Nieves (2025) generally yields results more compatible with those of Martini et al., some discrepancies remain. These differences could be precisely quantified once the revised approach of Ref. Sobczyk and Nieves (2025) is implemented in GENIE.
Figure 5 provides a specific example of a double differential cross section at fixed kinematics. To further illustrate the perfect correspondence between the theoretical calculations and the GENIE implementation in other configurations, we show in Fig. 6 the results for carbon at a different incident neutrino energy and for lepton scattering angle bins corresponding to those used in the T2K measurement Abe et al. (2020c) discussed in the next section.
III.2 Limitations of the model implementation
Whilst the implementation described here represents a complete inclusion of the content of the Martini et al. model within the GENIE event generator for a set of defined nuclear targets, there are limitations that are important to note. Some of these are related to approximations inherent in producing a fully exclusive simulations from an inclusive input model, which are common to many model implementations in neutrino interaction generators. Others are due to missing predictive power within the Martini et al. model itself. Some of these limitations can be mitigated with relatively modest extensions of the implementation.
Firstly, it is important to note that the hadron tensors for both 1p1h and npnh interactions currently only go up to in energy transfer and in momentum transfer. This is due to a choice made in the original implementation of the model Martini et al. (2009). In the current GENIE implementation, outside of this range, the differential cross section vanishes (like for the other models following the hadron tensor implementation method). Further tractable but non-trivial model development work would be required to overcome this limitation. In principle, the implementation can be artificially expanded by interpolating (“blending”) the cross sections predicted by the Martini et al. model with those from another inclusive model, such as SuSAv2 (as already done within the CRPA model implementation scheme in GENIE Dolan et al. (2022)). For the purposes of this implementation, as well as the comparisons with flux-averaged measurements we consider in this work, the current implementation is sufficient.
Importantly, the implementation scheme described in this document relies on a factorization scheme to generate hadron kinematics from an inclusive input model, like many implementations in neutrino interaction generators. The scheme followed is identical to that of Refs. Dolan et al. (2020, 2022) and is discussed further in Ref. Nikolakopoulos et al. (2023). This has only very small effects on the cross-section predictions shown within this paper since they have either no or only small dependence on the hadron kinematics predicted by GENIE.
We additionally note that, although the implementation features a 3p3h contribution to the npnh channel, this has been added as an additional “effective” 2p2h channel. This means that only two nucleons are produced at the vertex for a 3p3h interaction, relying solely on the existing GENIE hadron kinematics generator. This could be mitigated in a future extension to this implementation to produce a third final-state nucleon (via a new interaction channel within GENIE)666Separated 2p2h and 3p3h hadron tensors are already available.. Again, predictions in this article are either entirely unaffected or not very sensitive to alterations to the treatment of the hadronic system.
The current implementation relies on hadron tensors computed for three isoscalar nuclei – 12C, 16O and 40Ca. For nuclear targets other than those for which the tensors have been calculated, the total 1p1h cross section is scaled according to the difference in the number of target nucleons (i.e. depending on whether the incoming probe is a neutrino or an antineutrino). For example, for an 40Ar target, the total 1p1h cross section on 40Ca is scaled by a factor of 22/20 (18/20) for an incoming neutrino (antineutrino) probe. For npnh interactions, the scaling of the total cross section to different targets is based on atomic number, also following the scheme in Ref. Dolan et al. (2020). Note that a small additional scaling is applied, following the same scheme as for the SuSAv2 model Dolan et al. (2020) based on superscaling argument. This is the (square of) ratio of the Fermi momenta of the reference and target nuclei for 1p1h (npnh) interactions. Whilst not perfect, this overall scaling is expected to be a reasonable approximation away from the low energy transfer region (100 MeV).
More generally, the Martini et al. model internally does not differentiate between protons and neutrons when calculating cross sections, meaning that the impact of non-isoscalarity of target nuclei is not accounted for. Mitigating this limitation can be a challenging task. A first step could be assigning different Fermi momenta () to neutrons and protons within the model. However, this naive scaling does not take into account the modification of RPA suppression for asymmetric nuclei.
Finally, the Martini et al. model, and therefore the hadron tensors implemented in GENIE, do not take into account the impact of the nuclear removal energy. The consequence of this is that the lepton kinematics, in both GENIE and the Martini et al. model, are calculated assuming an interaction with unbound (but still moving) nucleons. On the other hand, GENIE does consider the removal energy when calculating hadron kinematics via the aforementioned factorization scheme. Whilst in this paper we do not attempt to modify the original inclusive predictions of the Martini et al. model, a possible mitigation for this limitation could be an extension of this implementation to add an effective modification to the energy transfer read from the hadron tensors to simulate the impact of removal energy. This strategy can be trivially adopted using routines already available within the GENIE generator.
Overall, the model implementation is a faithful reproduction of the Martini et al. model for outgoing lepton kinematics for carbon, oxygen or calcium nuclei. Scaling to different nuclei of similar atomic number is broadly reliable. The hadron kinematics predicted by the GENIE from this model implementation should be treated with cautious skepticism (as is the case also for other 1p1h and 2p2h GENIE model implementations).
IV Comparison with measurements
As discussed in Sec. II, the theoretical calculations by Martini et al. have already been compared in the past with several cross-section measurements. However, previous comparisons were performed in the absence of a unified generator implementation, so the treatment of effects such as final-state interactions (FSI) was not applied in a consistent way to all interaction channels. Here we present a few additional comparisons to illustrate the potential of the present GENIE implementation to make cross-section predictions for different neutrino beams, various nuclear targets and diverse final-state topologies, using a unified generator framework as enabled by this implementation.
To this end, we show the T2K flux-integrated charged-current double differential cross sections without pions in the final state, i.e. the “CC0” cross section, on carbon (Fig. 7(a)) and oxygen (Fig. 7(b)). The experimental T2K results are the regularized cross sections obtained via simultaneous analysis of the carbon and oxygen target samples, see Ref. Abe et al. (2020c) (and Fig. 9 in this publication) for more details. To process the simulated events, which include pion absorption contribution as implemented in GENIE, we use the NUISANCE framework Stowell et al. (2017) which facilitates direct measurement-model comparisons.
The Martini et al. model yields a reasonable agreement with the T2K measurement. The total for both carbon and oxygen distributions is for 58 degrees of freedom777A similar comparison with the so-called unregularized cross sections Abe et al. (2020c) yields for 58 degrees of freedom., which corresponds to a p-value of 0.05, indicating satisfactory agreement. This result can be compared with the values presented in Table V of Ref. Abe et al. (2020c) which presents the for nine different versions/tunes of the neutrino event generators888Note that some of these models have been updated since the measurement publication.. The obtained with the Martini et al. model is lower than that of five of these models and higher than that of four. It is useful to note that Table V of Ref. Abe et al. (2020c) emphasizes that the majority of the discrepancy between the models and the measurement is driven by the most forward bin, which is dominated by low energy transfer interactions. This is the region where suppression of the cross section from long-range correlations via RPA are most pronounced.
The Martini et al. model generally performs better than both the NEUT and NuWro Spectral Function (SF) models. This is expected, as these model implementations do not account for suppression of the cross section via RPA effects. The Martini et al. model also performs better than the GENIE SuSAv2 implementation and the RMF 1p1h+SuSAv2 prediction. In the first case, the SuSAv2 1p1h model implementation in GENIE does not include an explicit RPA suppression (but does include an effective suppression of the cross section at low energy transfer due to FSI, for more details see Dolan et al. (2020)), which results in an overprediction of the measurement in the most forward bins. In the case of the RMF model, the effect of FSI is included in the distortion of the final wavefunction from first principles, which causes a larger reduction in the cross section and thus a better agreement with the measurement. The Martini et al. model performs better than both of these models, indicating the probable need for an even larger suppression due to RPA effects. Finally, our implementation yields a lower than the GiBUU model tested in Ref. Abe et al. (2020c). This may be due to the fact that the LFG model used in GiBUU Buss et al. (2012) does not apply any RPA effects.
Conversely, the Martini et al. model yields a higher than the other GENIE, and NEUT Hayato and Pickering (2021) LFG+RPA-based models. In both cases, these implementations are based on the model of Nieves et al. Nieves et al. (2011) and, as discussed in Sec. III.1, have similar predictions for the inclusive cross section on carbon to those of the Martini et al. model (see Fig. 2 and Fig. 9 of Ref. Sobczyk and Nieves (2025)). A significant portion of the difference in the between the LFG-based models and the Martini et al. is therefore likely due to the treatment of 2p2h interactions. Whereas the updated treatment of multi-nucleon interactions discussed in Ref. Sobczyk and Nieves (2025) reduces the disagreement between the Martini et al. and the Nieves et al. models, the treatment used in the comparisons of Ref. Abe et al. (2020c) predicted a 2p2h cross section for the Nieves et al. model which is significantly lower than that of the Martini et al. (as shown in Fig. 2). It is additionally worth noting that the quoted values are for the combined carbon and oxygen measurement. A closer inspection of the individual carbon and oxygen inclusive predictions in the most forward bin reveals that the Nieves et al. and Martini et al. predictions are extremely close for carbon, but they differ slightly for oxygen. This implies that some missing ingredients in the detailed differences between the cross sections on carbon and oxygen affecting the forward region (for example concerning the nuclear removal energy) may be responsible for the higher .
Another example of model-measurement comparison is shown in Fig. 8 which presents the MicroBooNE flux integrated charged-current single differential cross sections without pions and with one proton in the final state – the so called “CC1p0” cross section – on argon. Compared to the previous T2K case, this measurement involves a different neutrino flux, a different nuclear target, and a different final-state topology. The measurement is that published by MicroBooNE in Ref. Abratenko et al. (2023), where the signal includes all -Ar scattering events with a final-state muon of momentum and exactly one proton with . Regarding the modeling, we recall that the implemented microscopic calculations are those of 40Ca, rescaled to 40Ar by following the same procedure discussed and justified in Ref. Martini et al. (2022), to which we refer the reader for details. Fig. 8 corresponds to Fig. 11 of Ref. Abratenko et al. (2023), which includes comparisons with other model predictions and the corresponding . In this case, the obtained with the Martini et al. model is for 18 degrees of freedom, which corresponds to a p-value of 0.994. Among the models considered, the Martini et al. model provides the lowest value for this particular measurement. This is likely due to the fact that the 2p2h cross section predicted by our implementation is significantly higher than that predicted by all of the other models considered in Fig. 11 of Ref. Abratenko et al. (2023), which are based on the previous version of the Nieves et al. model. Other studies (e.g. Ref. Filali et al. (2025)) have indeed shown that more exclusive MicroBooNE measurements indicate a preference for an enhanced 2p2h contribution on argon.
V Summary and Conclusion
The Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model has been implemented in the GENIE neutrino event generator. Even if this model allows a unified description of several channels (quasielastic, multinucleon emission and the coherent and incoherent single pion production) for both charged current and neutral current scattering, the present work concerns only the quasielastic and the multinucleon emission in the case of charged current (anti)neutrino scattering.
Basic validations have been performed and presented. All validation steps successfully passed.
Model predictions in GENIE for 12C, 16O and 40Ar have been compared with some available experimental measurements from T2K and MicroBooNE. Overall, a reasonable agreement has been observed.
This implementation of the Martini-Ericson-Chanfray-Marteau model in the widely used GENIE event generator is an important step towards eventual reduction of systematic uncertainties due to neutrino-nucleus cross section in a future combined analyses of data from the ongoing and next generation precision neutrino experiments.
Acknowledgments
MM thanks Magda Ericson, Guy Chanfray, Jacques Marteau and Jean Delorme for all the work that led to the development of the model implemented in the present manuscript. Magda Ericson and Guy Chanfray are also acknowledged by MM for the long and continuous collaboration. The authors would also like to acknowledge support and feedback on this work from the T2K and GENIE collaborations.
Appendix A Model predictions for 12C, 16O and 40Ca
In this Appendix we show the model predictions for the and charged-current total cross section as a function of the neutrino energy for three different nuclei which are of interest for the ongoing and future experiments: 12C (already presented and discussed in the main text), 16O and 40Ca (as a proxy for 40Ar). The corresponding comparisons are presented in Figs. 10 and 10 for and , respectively. A perfect agreement between the GENIE results and the original calculations is observed.

### v2

@@
-Implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model in the GENIE neutrino event generator
+Implementation of the Martini-Ericson-Chanfray-Marteau RPA-based neutrino and antineutrino cross-section model in the GENIE neutrino event generator
 Abstract
-We discuss the first implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model for quasielastic (1p1h) and multinucleon (2p2h and 3p3h) excitations in the widely used GENIE neutrino event generator.
-Validation steps are presented, in particular, through direct comparisons of GENIE cross section output with original calculations performed by the authors of the model.
-Predictions for 12C, 16O and 40Ar are compared with some available T2K and MicroBooNE experimental measurements showing a reasonable agreement.
+We discuss the first implementation of the Martini-Ericson-Chanfray-Marteau random phase approximation-based (anti)neutrino cross-section model for quasielastic (1p1h) and multinucleon (2p2h and 3p3h) excitations in the widely used GENIE neutrino event generator.
+Validation steps are presented, in particular, through direct comparisons of GENIE cross-section output with original calculations performed by the authors of the model.
+Predictions for 12C, 16O, and 40Ar are compared with some available T2K and MicroBooNE experimental measurements showing a reasonable agreement.
 I Introduction
@@
 (2018a) and DUNE Acciarri et al.
-(2015) will depend on the reduction to the percent level of the systematic errors.
+(2015a) will depend on the reduction to the percent level of the systematic errors.
 Presently, the most important source of uncertainties is represented by the neutrino-nucleus cross sections Katori and Martini (2018); Alvarez-Ruso et al.
@@
 Think, for example, of the neutrino energy reconstruction problem, the background subtraction for a peculiar experimental signal topology, or the modeling of the detector response.
-In this paper, we present the implementation of the theoretical RPA-based model of the Lyon group, described in detail in the article of Martini, Ericson, Chanfray and Marteau Martini et al.
+In this paper, we present the implementation of the theoretical random phase approximation (RPA) based model of the Lyon group, described in detail in the article of Martini et al.
+Martini et al.
 (2009) into the GENIE event generator Andreopoulos et al.
@@
 We then compare its predictions with different experimental measurements.
-This model allows a unified description of several scattering channels: quasielastic, multinucleon emission and the coherent and incoherent single pion production.
-Within this model, the debated MiniBooNE quasielastic-like cross section on carbon Aguilar-Arevalo et al.
+This model allows a unified description of several scattering channels: quasielastic (QE), multinucleon emission and the coherent and incoherent single pion production.
+Within this model, the debated MiniBooNE quasielasticlike cross section on carbon Aguilar-Arevalo et al.
 (2010) has been explained for the first time Martini et al.
-(2009, 2010) by stressing the crucial role of the multinucleon (also called 2p2h or npnh) excitations contribution which sizably increase the cross section.
+(2009, 2010) by stressing the crucial role of the multinucleon (also called 2p2h or npnh) excitation contributions which sizably increase the cross section.
 After this explanation, 2p2h excitations started to attract a lot of attention (for a review see, for example, Refs.
@@
 (2012); Ankowski et al.
-(2015), and consequently in the precise determination of the neutrino oscillation parameters.
-All widely-used event generators now include the multinucleon excitation channel, previously ignored before the MiniBooNE quasielastic-like results, into their packages.
-For a relatively recent state-of-the-art implementation in the different generators, see for example Refs.
+(2015) and, consequently, in the precise determination of the neutrino oscillation parameters.
+All widely-used event generators now include the multinucleon excitation channel, previously ignored before the MiniBooNE quasielasticlike results, into their packages.
+For a relatively recent state-of-the-art implementation in the different generators see, for example, Refs.
 Alvarez-Ruso et al.
@@
 (2021); Prasad et al.
+(2024); Mosel and Gallmeister (2024); Ankowski et al.
 (2025).
 This implementation effort is still ongoing, essentially for two reasons.
-Firstly, various theoretical models present some differences in their predictions for this channel, whose complexity – stemming from the intrinsic difficulty of nuclear physics – makes it hard to treat without approximations.
-Secondly, the experimental neutrino community is moving towards more exclusive measurements based on detected hadronic variables, whereas most of the existing models were originally developed for inclusive cross sections in terms of leptonic observables.
+First, various theoretical models present some differences in their predictions for this channel, whose complexity–stemming from the intrinsic difficulty of nuclear physics–makes it hard to treat without approximations.
+Second, the experimental neutrino community is moving toward more exclusive measurements based on detected hadronic variables, whereas most of the existing models were originally developed for inclusive cross sections in terms of leptonic observables.
+This interest in the hadronic variables is driven by the improved sensitivity to outgoing low-energy protons and neutrons of the upgraded T2K near detector ND280 Abe et al.
+(2019) as well as by the present Short-Baseline Neutrino Program at Fermilab Acciarri et al.
+(2015b) and the future DUNE experiment Acciarri et al.
+(2015a), both of which use liquid argon detectors with especially good hadronic sensitivity.
+The calorimetric energy estimator used there can also be problematically sensitive to the npnh effects (e.g.
+missing energy due to emitted neutrons or subthreshold protons).
 Up until now, the model of Ref.
 Martini et al.
-(2009) has not been included in any generators by the implementation of the corresponding theoretical hadron tensor, but only employing tables of shared cross sections at different kinematics via a re-weighting procedure Nguyen (2022).
+(2009) has not been included in any generators by the implementation of the corresponding theoretical hadron tensor, but only employing tables of shared cross sections at different kinematics via a reweighting procedure Nguyen (2022).
 The hadron tensor strategy is employed here for the first time for this model.
 For a detailed explanation of this strategy, see Ref.
+Schwehr et al.
+(2016) where it was developed as well as Ref.
 Dolan et al.
-(2020), where it is introduced in the context of the SuSAv2 model.
-Even though, in principle, all the channels described by this model can be implemented in the generators, in the present work we confine ourselves to considering only the quasielastic (also called 1p1h) and the multinucleon excitations channel (2p2h and 3p3h, collectively referred to as npnh), which have been the most investigated and employed in the neutrino analyses of accelerator-based experiments over the last fifteen years.
+(2020), where it is discussed in the context of the SuSAv2 model.
+Even though, in principle, all the channels described by this model can be implemented in the generators, in the present work we confine ourselves to considering only the quasielastic (also called 1p1h) and the multinucleon excitation channels (2p2h and 3p3h, collectively referred to as npnh), which have been the most investigated and employed in the neutrino analyses of accelerator-based experiments over the last fifteen years.
 For this purpose we follow the same protocol as the one adopted in Refs.
 Dolan et al.
-(2020) and Dolan et al.
-(2022) for the SuSAv2+MEC Megias et al.
-(2016) (quasielastic + 2p2h) and Ghent CRPA Pandey et al.
+(2020, 2022) for the SuSAv2+ meson-exchange current (MEC) Megias et al.
+(2016) (quasielastic + 2p2h) and Ghent continuum random-phase approximation (CRPA) Pandey et al.
 (2015) (giant resonances + quasielastic) models, respectively.
-The interest in implementing the Martini, Ericson, Chanfray, and Marteau model (hereafter referred to as the Martini et al.
+The interest in implementing the Martini-Ericson-Chanfray-and Marteau model (hereafter referred to as the Martini et al.
 model) relies on several aspects.
-First, as already mentioned, differences appear among the various models, and none is currently able to simultaneously describe all the available experimental measurements across the full neutrino energy range – from low (e.g., T2K) to high (e.g., MINERvA) energies – and across nuclear targets (most notably, 12C, 16O and 40Ar).
-In the absence of such a solution, the span of the predictions from plausible models already implemented inside generators provides some guidance about, for example, the size of systematic errors associated to different processes (e.g., see the heavy use of this approach in Ref.
+First, as already mentioned, differences appear among the various models, and none is currently able to simultaneously describe all the available experimental measurements across the full neutrino energy range–from low (e.g., T2K) to high (e.g., MINERvA) energies–and across nuclear targets (most notably, 12C, 16O, and 40Ar).
+In the absence of such a solution, the span of the predictions from plausible models already implemented inside generators provides some guidance about, for example, the size of systematic errors associated with different processes (e.g., see the heavy use of this approach in Ref.
 Wilkinson et al.
@@
 Second, in the Martini et al.
-model the different 2p2h contributions – namely the ones related to i) nucleon-nucleon correlations, ii) meson exchange currents (MEC), and iii) their interference – as well as the 3p3h excitations, are separately calculated and available as distinct components of the hadron tensor.
+model, the different 2p2h contributions–namely the ones related to (i) nucleon-nucleon correlations, (ii) meson exchange currents, and (iii) their interference–as well as the 3p3h excitations, are separately calculated and available as distinct components of the hadron tensor.
 This allows for easier comparison between 2p2h models and, overall, could help avoid the risk of double counting (e.g., nucleon-nucleon correlations contributions) when combining different channels taken from different models (e.g., using a spectral function model, like that in Ref.
 Benhar and Meloni (2007), for the quasielastic channel and Fermi gas-based models for 2p2h interactions).
-Finally, calculations for different atomic nuclei are available, enabling the modeling of cross sections for carbon, water and argon targets – the main targets of present and future accelerator-based neutrino experiments.
+Finally, calculations for 12C, 16O and 40Ca are available, enabling the modeling of cross sections for carbon, water, and argon targets–the main targets of present and future accelerator-based neutrino experiments.
+Even if the Martini et al.
+model allows also the treatment of neutral-current (NC) neutrino cross sections, and the nuclear responses are available also for the electron scattering, in the present work we focus exclusively on the GENIE implementation of the neutrino and antineutrino (1p1h and npnh) charged-current (CC) cross sections.
 The article is organized as follows.
@@
 (2009), to which we refer for all the details.
-It is based on nuclear response functions treated in the random phase approximation (RPA) on top of a local Fermi gas (LFG) calculation.
+It is based on nuclear response functions treated in the RPA on top of a local Fermi gas (LFG) calculation.
 This formalism belongs to a long tradition of investigations: first on electron scattering and pionic modes in nuclei (see, for example, Refs.
@@
 (1993); Alberico et al.
-(1998)), and later on neutrino scattering Delorme and Ericson (1985); Marteau (1999); Marteau et al.
+(1998)) and, later on, neutrino scattering Delorme and Ericson (1985); Marteau (1999); Marteau et al.
 (2000).
@@
 One of its strengths is that it allows a unified description of several scattering channels: quasielastic, multinucleon emission and the coherent and incoherent single pion production.
-As already mentioned, within this model, the debated MiniBooNE quasielastic-like cross section on carbon Aguilar-Arevalo et al.
+As already mentioned, within this model, the debated MiniBooNE quasielasticlike cross section on carbon Aguilar-Arevalo et al.
 (2010) was explained for the first time Martini et al.
@@
 The Martini et al.
-model has successfully reproduced many of the features of several muon-neutrino flux-integrated double differential cross-section measurements in the past: the MiniBooNE Charged Current Quasielastic (CCQE) - like neutrino Martini et al.
+model has successfully reproduced many of the features of several muon-neutrino flux-integrated double differential cross-section measurements in the past: the MiniBooNE charged-current quasielasticlike neutrino Martini et al.
 (2011) and antineutrino Martini and Ericson (2013) measurements, as well as their combination Ericson and Martini (2015); the MiniBooNE CC1 and the T2K CC inclusive measurements Martini and Ericson (2014); Martini et al.
@@
 (2024a) cross sections on argon.
-The counter-intuitive dominance of the differential cross sections over the ones in specific kinematical conditions – not only for quasielastic, but also for multinucleon excitations and single pion production – has been explained with simple arguments in Ref.
+The counterintuitive dominance of the differential cross sections over the ones in specific kinematical conditions–not only for quasielastic, but also for multinucleon excitations and single pion production–has been explained with simple arguments in Ref.
 Martini et al.
@@
 (2017, 2018b, 2020b, 2023b, 2025).
-In these experimental papers, the model is often referred to as “Martini et al.” or “Lyon”.
-The main reason for considering it is its larger cross-section predictions in the 2p2h sector (especially for neutrinos, and to a lesser extent for antineutrinos) compared to those of Ref.
+In these experimental papers, the model is often referred to as “Martini et al.” or “Lyon.” The main reason for considering it is its larger cross-section predictions in the 2p2h sector (especially for neutrinos and, to a lesser extent, for antineutrinos) compared to those of Ref.
 Nieves et al.
-(2011), a similar theoretical approach, often referred to as “Nieves et al.” or “Valencia”, which is already implemented in many generators.
+(2011), a similar theoretical approach, often referred to as “Nieves et al.” or “Valencia,” which is already implemented in many generators.
 A very recent reexamination Sobczyk and Nieves (2025) of the Nieves et al.
 approach now leads to new results that are closer to that of Martini et al.
+Before concluding this section, for clarity and to make the work as self-contained as possible (without entering in too much detail already published elsewhere) we remind the reader of some general aspects of the neutrino-nucleus cross sections and of the formalism employed by Martini et al.
+in their calculations.
+The double differential cross section for the charged-current reaction is given by
+Here is the differential solid angle in the direction specified by the charged lepton momentum in the laboratory frame, is the energy transferred to the nucleus, the zero component of the four-momentum transfer , with and , being the initial and final lepton four momenta.
+In Eq.
+(1), is the weak coupling constant, is the Cabibbo angle.
+is the leptonic tensor
+and is the hadronic tensor, a combination of nucleon form factors (that contain the information about the nucleon properties), and nuclear response functions (that contain the information about the nuclear dynamics).
+The different components of the hadronic tensor can be combined allowing a reformulation of Eq.
+(1) in terms of projections with respect to the momentum transfer direction.
+The charged-current cross section is a linear combination of five contributions
+where the kinematical factors come from the contraction with the leptonic tensor and the plus (minus) sign applies to neutrinos (antineutrinos).
+We do not provide here the explicit expressions for the kinematical factors and for the nucleon form factors and nuclear responses entering in the hadronic tensors, as they can be found in Appendixes A and B of Ref.
+Martini et al.
+(2009).111There is a typo in the expression of in Eq.
+(A4) of Ref.
+Martini et al.
+(2009): the factor 8 should be replaced by the factor 16.
+The hadronic tensor can be decomposed in terms of the following scattering channels:
+that are available as subtables of tensor elements. It can also be decomposed in terms of the different channels of particle-hole (nucleon or Delta ) excitations
+which are particularly useful to classify the type of 2p2h contributions.
+It is important to remember that Eqs.
+(1) and (3) are totally general and, according to the decompositions of Eq.
+(4), apply to different excitations channels: 1p1h quasielastic, 2p2h, 3p3h (the three channels considered in the present work), and 1p1h 1 production, coherent production (the two other channels of the Martini et al.
+model, whose Monte Carlo implementation is left for future works).
 Since the present work focuses mainly on multinucleon excitations, we recall that in the model of Ref.
@@
 Martini et al.
-(2009)); the -MEC contribution (denoted in Ref.Martini et al.
-(2009)); and the NN correlations–-MEC interference, (denoted N in Ref.
+(2009)); the -MEC contribution (denoted in Ref Martini et al.
+(2009)); and the NN correlations–-MEC interference (denoted N in Ref.
 Martini et al.
 (2009)).
+These contributions are separately available as subtable of the hadron tensor , according to the decompositions of Eq.
+(5) , which holds for the different scattering channels.
 We also recall that, as in all the works of the Lyon group starting from the end of Ref.
@@
 (2000).
-This so-called “new” evaluation is based, for the NN and N parts, on the microscopic calculations of Ref.
+This so-called new evaluation is based, for the NN and N parts, on the microscopic calculations of Ref.
 Alberico et al.
 (1984), aimed at determining the 2p2h contribution to the isospin spin-transverse response.
-For the part, the contributions related to the non-pionic decay of the inside the nuclear medium – leading to 2p2h and 3p3h excitations – are taken from Ref.
-Oset and Salcedo (1987), while the terms not reducible to a modifications of the width are taken from Ref.
+For the part, the contributions related to the nonpionic decay of the inside the nuclear medium–leading to 2p2h and 3p3h excitations–are taken from Ref.
+Oset and Salcedo (1987), while the terms not reducible to modifications of the width are taken from Ref.
 Delorme and Guichon (1991), which exploits the calculations of Ref.
 Shimizu and Faessler (1980).
-For further details, see Sec.
-IV C of Ref.
-Martini et al.
-(2009) and Sec.
-II of Ref.
-Martini et al.
-(2010).
-We also recall that all the form factors entering the hadron tensor coincide with those of Ref.
+Further details are given in Appendix B.
+We also recall that, for the purpose of validating the present implementation against the previously published theoretical results, all the form factors entering the hadron tensor coincide with those of Ref.
 Martini et al.
 (2009), with exactly the same parameters.
-For example, for the axial form factor, a standard dipole parameterization is adopted, with and .
+For example, for the axial form factor, a standard dipole parametrization is adopted, with and the value222We remind the reader that the PDG 2024 Navas et al.
+(2024) value of the axial coupling constant is slightly different, being .
+The effects of this difference with respect to our adopted value are expected to be small.
+They could be quantified in a future work.
+of .
 III Implementation in GENIE and validation
@@
 Dolan et al.
-(2020) and Dolan et al.
-(2022).
+(2020, 2022).
 This approach relies on the fact that the differential cross section can be expressed as the product of kinematic factors and the contraction of a generic lepton tensor with a model-specific hadron tensor.
@@
 Schwehr et al.
-(2016), Dolan et al.
-(2020) and Dolan et al.
-(2022).
-These tables are provided for the 1p1h and npnh channels for carbon, oxygen and calcium targets.
+(2016); Dolan et al.
+(2020, 2022).
+These tables are provided for the 1p1h and npnh channels for carbon, oxygen, and calcium targets.
 The npnh contribution includes both 2p2h and 3p3h components, and separate hadron tensors are employed for each.
@@
 III.2.
-The input hadron tensors are defined over an energy transfer varying in the range , and a momentum transfer varying in the range , as in Ref.
+The input hadron tensors are defined over an energy transfer varying in the range 5 and a momentum transfer varying in the range 1 , as in Ref.
 Martini et al.
@@
 Dolan et al.
-(2020) and Dolan et al.
-(2022), similar to the one described in Ref.
+(2020, 2022), similar to the one described in Ref.
 Schwehr et al.
@@
 2 and 2, respectively (and in Figs.
-10 and 10 of Appendix References for 16O and 40Ca).
+10 and 10 of Appendix A for 16O and 40Ca).
 These plots demonstrate perfect agreement between the original calculations (shaded bands) and the GENIE predictions for the different contributions to the cross section: “1p1h” (dashed), “npnh” (dotted) and their sum (solid).
-The predictions of other 1p1h and 2p2h models available in GENIE (“Nieves et al.” and “SuSAv2”) are also shown for reference.
+The predictions of other 1p1h and 2p2h models available in GENIE (Nieves et al.
+and SuSAv2) are also shown for reference.
 The corresponding GENIE tunes used are G18_10a_00_000 for the Nieves et al.
 model and G21_11a_00_000 for SuSAv2.
-As already illustrated in other articles (see for example Fig.
+As already illustrated in other articles (see, for example, Fig.
 9 of Ref.
 Abe et al.
-(2023a)) the npnh (as well as the 1p1h) contributions estimated by these three models are different, and their relative importance with respect to 1p1h varies between neutrinos and antineutrinos, as discussed in Refs.
+(2023a)), the npnh (as well as the 1p1h) contributions estimated by these three models are different, and their relative importance with respect to 1p1h varies between neutrinos and antineutrinos, as discussed in Refs.
 Martini et al.
@@
 2 of Dolan et al.
-(2020) respectively.
+(2020), respectively.
 In Fig.
@@
 where denotes the nucleon mass.
-The spread around this curve is due to Fermi motion111The fact that the quasielastic region is not strictly limited by the two quasielastic bands , typical of a global relativistic Fermi gas-based model with Fermi momentum , is due to the local density approximation and to the presence of Bessel functions in the nuclear response definitions (see Appendix 2 of Ref.
+The spread around this curve is due to Fermi motion333The fact that the quasielastic region is not strictly limited by the two quasielastic bands , typical of a global relativistic Fermi gas-based model with Fermi momentum , is due to the local density approximation and to the presence of Bessel functions in the nuclear response definitions (see Appendix 2 of Ref.
 Martini et al.
@@
 Dolan et al.
-(2020) – arising from nucleon-nucleon correlations, and a second one above the quasielastic region, extending from the dip to the region, due to -MEC contributions222These two branches are more distinguishable looking directly to the phase space of the nuclear response function, rather than to the flux-integrated cross section.
+(2020)–arising from nucleon-nucleon correlations, and a second one above the quasielastic region, extending from the dip to the region, due to -MEC contributions444These two branches are more distinguishable looking directly to the phase space of the nuclear response function, rather than to the flux-integrated cross section.
 See for example Fig.
@@
 model differs from those of Nieves et al.
-and SuSAv2, though it shares features with both: the presence of two branches, as in Nieves et al., and the accumulation of cross section strength in the dip region, as in SuSAv2.
+and SuSAv2, though it shares features with both: the presence of two branches, as in Nieves et al., and the accumulation of cross-section strength in the dip region, as in SuSAv2.
 As a validation of the implementation, and to facilitate comparisons with other models, in Fig.
@@
 Assuming azimuthal symmetry, the double differential cross section is related by .
-for both neutrino and antineutrino interactions on oxygen, at fixed incident energy and lepton scattering angle444GENIE does not allow the user to explicitly fix the scattering angle.
-Events with were selected to approximate a scattering angle of ., as a function of the energy transfer .
+for both neutrino and antineutrino interactions on oxygen, at fixed incident energy and lepton scattering angle,666GENIE does not allow the user to explicitly fix the scattering angle.
+Events with were selected to approximate a scattering angle of .
+as a function of the energy transfer .
 This represents the most stringent validation test, as the cross sections are not integrated over any variable, in contrast to the total or flux-integrated differential cross sections discussed earlier.
-A perfect agreement is observed between the theoretical calculations (solid lines) and the GENIE implementation (histograms) for both 1p1h and npnh channels, and for neutrino as well as antineutrino interactions555In the 1p1h case, the small fluctuations at low discussed in footnote 1 can also be seen..
-Fig.
-5 can be directly compared with Fig.
+A perfect agreement is observed between the theoretical calculations (solid lines) and the GENIE implementation (histograms) for both 1p1h and npnh channels, and for neutrino as well as antineutrino interactions.777In the 1p1h case, the small fluctuations at low discussed in footnote 3 can also be seen.
+Figure 5 can be directly compared with Fig.
 9 of Ref.
@@
 5 is slightly more peaked and shifted to lower by approximately .
-This is probably due to the absence of removal energy in the calculations of Martini et al.
-unlike the case of Nieves et al.
+This is probably due to the absence of removal energy in the calculations of Martini et al., unlike the case of Nieves et al.
 These small differences are often washed out in flux-integrated double differential distributions (see, e.g., Fig.
@@
 III.2 Limitations of the model implementation
-Whilst the implementation described here represents a complete inclusion of the content of the Martini et al.
-model within the GENIE event generator for a set of defined nuclear targets, there are limitations that are important to note.
+While the implementation described here represents an inclusion of the 1p1h and npnh channels of the Martini et al.
+model within the GENIE event generator for a set of defined nuclear targets for charged-current neutrino and antineutrino scattering, there are limitations that are important to note.
 Some of these are related to approximations inherent in producing a fully exclusive simulations from an inclusive input model, which are common to many model implementations in neutrino interaction generators.
@@
 Some of these limitations can be mitigated with relatively modest extensions of the implementation.
-Firstly, it is important to note that the hadron tensors for both 1p1h and npnh interactions currently only go up to in energy transfer and in momentum transfer.
+First, it is important to note that the hadron tensors for both 1p1h and npnh interactions currently only go up to in energy transfer and in momentum transfer.
 This is due to a choice made in the original implementation of the model Martini et al.
@@
 In the current GENIE implementation, outside of this range, the differential cross section vanishes (like for the other models following the hadron tensor implementation method).
-Further tractable but non-trivial model development work would be required to overcome this limitation.
+Further tractable but nontrivial model development work would be required to overcome this limitation.
 In principle, the implementation can be artificially expanded by interpolating (“blending”) the cross sections predicted by the Martini et al.
@@
 For the purposes of this implementation, as well as the comparisons with flux-averaged measurements we consider in this work, the current implementation is sufficient.
+We remind the reader also that, even if the tensors start from , giant resonance peaks that may appear at low transferred momenta for are not included in the formalism of Martini et al.
+at difference of the CRPA Pandey et al.
+(2015); Dolan et al.
+(2022) approach, particularly suited for this low-energy low-momentum region.
+However, in this region, the RPA Martini et al.
+calculations (which do not include giant resonance peaks) may represent a reasonable average of the CRPA results Pandey et al.
+(2015), as shown in Fig.
+2 of Ref.
+Martini et al.
+(2016) and discussed in that article.
 Importantly, the implementation scheme described in this document relies on a factorization scheme to generate hadron kinematics from an inclusive input model, like many implementations in neutrino interaction generators.
@@
 This means that only two nucleons are produced at the vertex for a 3p3h interaction, relying solely on the existing GENIE hadron kinematics generator.
-This could be mitigated in a future extension to this implementation to produce a third final-state nucleon (via a new interaction channel within GENIE)666Separated 2p2h and 3p3h hadron tensors are already available..
+This could be mitigated in a future extension to this implementation to produce a third final-state nucleon (via a new interaction channel within GENIE)888We note that separate 2p2h and 3p3h hadron tensors are available if the user prefers to only simulate one of the two npnh contributions, or a specific sub-contribution within the 2p2h channel as described in the final paragraph of Sec.
+II.
+This option can be configured by editing the config/MartiniMECHadronTensorModel.xml configuration file..
+This was not implemented at the time of writing this article as the current implementation uses the same routine to process the combined npnh (i.e.
+2p2h+3p3h) tensors.
+Separating the two types of interactions would have required creating a new routine for the 3p3h channel.
+Since the current model predictions are inclusive, we leave this extension for future work.
 Again, predictions in this article are either entirely unaffected or not very sensitive to alterations to the treatment of the hadronic system.
@@
 For example, for an 40Ar target, the total 1p1h cross section on 40Ca is scaled by a factor of 22/20 (18/20) for an incoming neutrino (antineutrino) probe.
-For npnh interactions, the scaling of the total cross section to different targets is based on atomic number, also following the scheme in Ref.
+For npnh interactions, the total cross section is scaled with the mass number, also following the scheme in Ref.
 Dolan et al.
+(2020), implying no scaling between the 40Ca and 40Ar npnh cross sections.
+Note that a small additional scaling is applied, following the same scheme as for the SuSAv2 model Dolan et al.
+(2020) based on superscaling arguments.
+This is the (square of the) ratio of the Fermi momenta of the reference and target nuclei for 1p1h (npnh) interactions.999The Fermi momentum used for this calculation are taken from Ref.
+Maieron et al.
 (2002).
-Note that a small additional scaling is applied, following the same scheme as for the SuSAv2 model Dolan et al.
-(2020) based on superscaling argument.
-This is the (square of) ratio of the Fermi momenta of the reference and target nuclei for 1p1h (npnh) interactions.
-Whilst not perfect, this overall scaling is expected to be a reasonable approximation away from the low energy transfer region (100 MeV).
+It is assumed to be the same for protons and neutrons.
+For nuclei not in the table the Fermi momentum of the nucleus with the closest proton number is used.
+While not perfect, this overall scaling is expected to be a reasonable approximation away from the low-energy transfer region (100 MeV).
+Scaling to heavier nuclear targets occasionally employed in neutrino scattering measurements, such as iron or lead, could be improved with a future addition of a hadron tensor table for a heavier target.
 More generally, the Martini et al.
-model internally does not differentiate between protons and neutrons when calculating cross sections, meaning that the impact of non-isoscalarity of target nuclei is not accounted for.
+model internally does not differentiate between protons and neutrons when calculating cross sections, meaning that the impact of nonisoscalarity of target nuclei is not accounted for.
 Mitigating this limitation can be a challenging task.
 A first step could be assigning different Fermi momenta () to neutrons and protons within the model.
-However, this naive scaling does not take into account the modification of RPA suppression for asymmetric nuclei.
+However, this naive scaling does not take into account the RPA modification for asymmetric nuclei Alberico et al.
+(1989), a challenge that, to our knowledge, is shared with the other RPA-based models at present, theoretical calculations of asymmetric relativistic Fermi gas for quasielastic lepton-nucleus scattering being only relatively recent Barbaro et al.
+(2018).
+It is also important to stress that the first microscopic calculation of two-particle-two-hole meson-exchange currents in 40Ar discussing also asymmetric scaling properties for neutrino and electron scattering appeared only very recently Martinez-Consentino et al.
+(2025).
 Finally, the Martini et al.
@@
 On the other hand, GENIE does consider the removal energy when calculating hadron kinematics via the aforementioned factorization scheme.
-Whilst in this paper we do not attempt to modify the original inclusive predictions of the Martini et al.
+While in this paper we do not attempt to modify the original inclusive predictions of the Martini et al.
 model, a possible mitigation for this limitation could be an extension of this implementation to add an effective modification to the energy transfer read from the hadron tensors to simulate the impact of removal energy.
@@
 Scaling to different nuclei of similar atomic number is broadly reliable.
-The hadron kinematics predicted by the GENIE from this model implementation should be treated with cautious skepticism (as is the case also for other 1p1h and 2p2h GENIE model implementations).
+The hadron kinematics predicted by GENIE from this model implementation should be treated with cautious skepticism (as is the case also for other 1p1h and 2p2h GENIE model implementations).
 IV Comparison with measurements
@@
 have already been compared in the past with several cross-section measurements.
-However, previous comparisons were performed in the absence of a unified generator implementation, so the treatment of effects such as final-state interactions (FSI) was not applied in a consistent way to all interaction channels.
+However, previous comparisons were performed in the absence of a unified generator implementation, so the treatment of effects such as final-state interactions (FSIs) was not applied in a consistent way to all interaction channels.
 Here we present a few additional comparisons to illustrate the potential of the present GENIE implementation to make cross-section predictions for different neutrino beams, various nuclear targets and diverse final-state topologies, using a unified generator framework as enabled by this implementation.
-To this end, we show the T2K flux-integrated charged-current double differential cross sections without pions in the final state, i.e.
+Throughout this section, we compare to two experimental measurements from the T2K and MicroBooNE Collaborations, both of which report the cross section for the so-called CC topology.
+This topology contains contributions from the 1p1h and npnh channels, which are simulated with the Martini et al.
+model implementation described in this article.
+There is also a subleading contribution from interactions where a pion was produced (usually via a resonant interaction) and then absorbed via FSI.
+To model this contribution, we use the standard GENIE prescription (Ref.
+Alvarez-Ruso et al.
+(2021)) which employs the Berger-Sehgal model (Ref.
+Berger and Sehgal (2007)) with updated form factors from Graczyk and Sobczyk (Ref.
+Graczyk and Sobczyk (2008)) and includes a variety of resonances up to invariant hadronic mass GeV, but neglecting interferences between them.
+Final-state interactions are simulated via the GENIE intranuclear cascade using the hA2018Intranuke routine (for details, see Ref.
+Dytman et al.
+(2021)).
+FSIs will impact the predictions in several ways.
+Although the FSI model in GENIE does not modify the total inclusive cross section, it will affect the hadrons that are used in defining the signal topologies.
+For both T2K and MicroBooNE, FSI impacts the samples mainly via pion absorption, which contributes to the measured signal sample.
+In addition to this, the MicroBooNE samples require at least one proton with momentum above , and the final-state proton kinematics that determine the sample composition will also be impacted by the choice of the FSI model (its effect is notably to decelerate protons and thus reduce the total cross section).
+It is also worth noting that the MicroBooNE samples also allow final-state charged pions with momenta below , which can also be affected by the choice of resonant pion production and FSI models.
+Finally, we note that FSI can also impact both T2K and MicroBooNE samples via pion production from nucleon interactions with the residual nucleus, but the contribution of these processes in this energy range is negligible for the purposes of this discussion.
+We begin by showing the T2K flux-integrated charged-current double differential cross sections without pions in the final state, i.e.
 the “CC0” cross section, on carbon (Fig.
@@
 model yields a reasonable agreement with the T2K measurement.
-The total for both carbon and oxygen distributions is for 58 degrees of freedom777A similar comparison with the so-called unregularized cross sections Abe et al.
-(2020c) yields for 58 degrees of freedom., which corresponds to a p-value of 0.05, indicating satisfactory agreement.
+The total for both carbon and oxygen distributions is for 58 degrees of freedom,101010A similar comparison with the so-called unregularized cross sections Abe et al.
+(2020c) yields for 58 degrees of freedom.
+which corresponds to a p value of 0.05, indicating satisfactory agreement.
 This result can be compared with the values presented in Table V of Ref.
 Abe et al.
-(2020c) which presents the for nine different versions/tunes of the neutrino event generators888Note that some of these models have been updated since the measurement publication..
+(2020c) which presents the for nine different versions/tunes of the neutrino event generators.111111Note that some of these models have been updated since the measurement publication.
 The obtained with the Martini et al.
@@
 Abe et al.
-(2020c) emphasizes that the majority of the discrepancy between the models and the measurement is driven by the most forward bin, which is dominated by low energy transfer interactions.
-This is the region where suppression of the cross section from long-range correlations via RPA are most pronounced.
+(2020c) emphasizes that the majority of the discrepancy between the models and the measurement is driven by the most forward bin, which is dominated by low-energy transfer interactions.
+This is the region where suppression of the cross section from long-range correlations via RPA is most pronounced.
 The Martini et al.
-model generally performs better than both the NEUT and NuWro Spectral Function (SF) models.
+model generally performs better than both the NEUT and NuWro spectral function models.
 This is expected, as these model implementations do not account for suppression of the cross section via RPA effects.
@@
 model also performs better than the GENIE SuSAv2 implementation and the RMF 1p1h+SuSAv2 prediction.
-In the first case, the SuSAv2 1p1h model implementation in GENIE does not include an explicit RPA suppression (but does include an effective suppression of the cross section at low energy transfer due to FSI, for more details see Dolan et al.
-(2020)), which results in an overprediction of the measurement in the most forward bins.
+In the first case, the SuSAv2 1p1h model implementation in GENIE does not include an explicit RPA suppression (but does include an effective suppression of the cross section at low-energy transfer due to FSI; for more details, see Dolan et al.
+(2020)), which results in an overprediction of the measurement in the most forward cos bins.
 In the case of the RMF model, the effect of FSI is included in the distortion of the final wave function from first principles, which causes a larger reduction in the cross section and thus a better agreement with the measurement.
@@
 Conversely, the Martini et al.
-model yields a higher than the other GENIE, and NEUT Hayato and Pickering (2021) LFG+RPA-based models.
+model yields a higher than the other GENIE and NEUT Hayato and Pickering (2021) LFG+RPA-based models.
 In both cases, these implementations are based on the model of Nieves et al.
@@
 III.1, have similar predictions for the inclusive cross section on carbon to those of the Martini et al.
-model (see Fig.
-2 and Fig.
-9 of Ref.
+model (see Figs.
+2 and 9 of Ref.
 Sobczyk and Nieves (2025)).
 A significant portion of the difference in the between the LFG-based models and the Martini et al.
-is therefore likely due to the treatment of 2p2h interactions.
-Whereas the updated treatment of multi-nucleon interactions discussed in Ref.
+model is therefore likely due to the treatment of 2p2h interactions.
+Whereas the updated treatment of multinucleon interactions discussed in Ref.
 Sobczyk and Nieves (2025) reduces the disagreement between the Martini et al.
@@
 Another example of model-measurement comparison is shown in Fig.
-8 which presents the MicroBooNE flux integrated charged-current single differential cross sections without pions and with one proton in the final state – the so called “CC1p0” cross section – on argon.
+8, which presents the MicroBooNE flux-integrated charged-current single differential cross sections without pions and with one proton in the final state–the so called CC1p0 cross section–on argon.
 Compared to the previous T2K case, this measurement involves a different neutrino flux, a different nuclear target, and a different final-state topology.
@@
 (2022), to which we refer the reader for details.
-Fig.
-8 corresponds to Fig.
+Figure 8 corresponds to Fig.
 11 of Ref.
@@
 In this case, the obtained with the Martini et al.
-model is for 18 degrees of freedom, which corresponds to a p-value of 0.994.
+model is for 18 degrees of freedom, which corresponds to a p value of 0.994.
 Among the models considered, the Martini et al.
@@
 Abratenko et al.
-(2023), which are based on the previous version of the Nieves et al.
-model.
-Other studies (e.g.
-Ref.
+(2023), which are all based on the previous version of the Nieves et al.
+model but GiBUU.121212We notice that in Fig.
+11 of Ref.
+Abratenko et al.
+(2023) GiBUU provides the lowest () with respect to the other models considered in that figure.
+This reinforces our hypothesis that the reason is related to the 2p2h contribution, the GiBUU one being larger than the one of the other Monte Carlos, all based on the previous version of the Nieves et al.
+model, and slightly lower than the predictions of Martini et al.
+for argon, as one can appreciate by comparing Fig.
+8 of Ref.
+Mosel and Gallmeister (2024) with Fig.
+5 of Ref.
+Martini et al.
+(2022).
+Other studies (e.g., Ref.
 Filali et al.
@@
 The Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model has been implemented in the GENIE neutrino event generator.
-Even if this model allows a unified description of several channels (quasielastic, multinucleon emission and the coherent and incoherent single pion production) for both charged current and neutral current scattering, the present work concerns only the quasielastic and the multinucleon emission in the case of charged current (anti)neutrino scattering.
+Even if this model allows a unified description of several channels (quasielastic, multinucleon emission and the coherent and incoherent single pion production) for both charged- and neutral-current scattering, the present work concerns only the quasielastic and the multinucleon emission in the case of charged-current (anti)neutrino scattering.
+The implementation of single pion production as well as of electromagnetic and neutral-current scattering is left for future works.
 Basic validations have been performed and presented. All validation steps successfully passed.
-Model predictions in GENIE for 12C, 16O and 40Ar have been compared with some available experimental measurements from T2K and MicroBooNE. Overall, a reasonable agreement has been observed.
-This implementation of the Martini-Ericson-Chanfray-Marteau model in the widely used GENIE event generator is an important step towards eventual reduction of systematic uncertainties due to neutrino-nucleus cross section in a future combined analyses of data from the ongoing and next generation precision neutrino experiments.
+Model predictions in GENIE for 12C, 16O, and 40Ar have been compared with some available experimental measurements from T2K and MicroBooNE. Overall, a reasonable agreement has been observed.
+This implementation of the Martini-Ericson-Chanfray-Marteau model in the widely used GENIE event generator is an important step toward eventual reduction of systematic uncertainties due to neutrino-nucleus cross section in future combined analyses of data from the ongoing and next generation precision neutrino experiments.
 Acknowledgments
-MM thanks Magda Ericson, Guy Chanfray, Jacques Marteau and Jean Delorme for all the work that led to the development of the model implemented in the present manuscript.
-Magda Ericson and Guy Chanfray are also acknowledged by MM for the long and continuous collaboration.
+M.
+M.
+thanks Magda Ericson, Guy Chanfray, Jacques Marteau and Jean Delorme for all the work that led to the development of the model implemented in the present manuscript.
+Magda Ericson and Guy Chanfray are also acknowledged by M.
+M.
+for the long and continuous collaboration.
 The authors would also like to acknowledge support and feedback on this work from the T2K and GENIE Collaborations.
 Appendix A Model predictions for 12C, 16O and 40Ca
-In this Appendix we show the model predictions for the and charged-current total cross section as a function of the neutrino energy for three different nuclei which are of interest for the ongoing and future experiments: 12C (already presented and discussed in the main text), 16O and 40Ca (as a proxy for 40Ar).
+In this appendix we show the model predictions for the and charged-current total cross section as a function of the neutrino energy for three different nuclei that are of interest for the ongoing and future experiments: 12C (already presented and discussed in the main text), 16O, and 40Ca (as a proxy for 40Ar).
 The corresponding comparisons are presented in Figs.
@@
 A perfect agreement between the GENIE results and the original calculations is observed.
+Appendix B A brief summary of the npnh treatment in the Martini-Ericson-Chanfray-Marteau model
+In this appendix we describe how npnh excitations are treated in the Martini et al.
+model by following an historical perspective which summarizes Sec.
+IV C of Ref.
+Martini et al.
+(2009) and Sec.
+II of Ref.
+Martini et al.
+(2010).
+There are several sources of npnh excitations.
+One of these sources arises from the modification of the width in the nuclear medium.
+This effect was introduced and parametrized by Oset and Salcedo Oset and Salcedo (1987), which considered the two-body (2p2h) [Fig.
+11(d)] and the three-body (3p3h) channels in the case of real pion or photon absorption.
+Martini et al.
+have used their parametrization of the modified width, although the kinematics of neutrino interaction is different since we are in the spacelike region, which could be a source of uncertainty.
+For the other terms not reducible to a modification of a width, of which some examples are shown in Fig.
+11, they have first used, as in the papers of Marteau et al.
+Marteau (1999); Marteau et al.
+(2000), a parametrization of Delorme and Guichon Delorme and Guichon (1989) (called “old” in Sec.
+IV C of Ref.
+Martini et al.
+(2009)).
+They exploited a calculation by Shimizu and Faessler Shimizu and Faessler (1980) of the absorptive part of the p-wave pion-nucleus optical potential at threshold.
+It is known that the absorption mechanism of pions is a two-nucleon one, which means that in the final state two nucleons are ejected.
+This is a 2p2h excitation.
+In Ref.
+Shimizu and Faessler (1980), the absorption is described by three types of terms (see Fig.
+11).
+The first one [Fig.
+11(a) with pion lines replacing the weak current ones as in the next diagrams] arises from the correlations, essentially from the tensor correlations.
+Another one involves a excitation [Fig.
+11(c)].
+The third one is an interference between the nucleon correlation and terms ( interference) as in Fig.
+11(b).
+The pion absorption calculation of Shimizu and Faessler Shimizu and Faessler (1980) is performed for threshold pion, i.e., for a vanishing three-momentum and an energy , which does not correspond to the neutrino situation.
+Delorme and Guichon Delorme and Guichon (1989) have then introduced in each absorption graph the corresponding energy dependence to obtain the bare 2p2h responses to be inserted in the RPA chain needed for neutrino interaction.
+The explicit expressions of the corresponding propagators are given in Appendix B 1 of Ref.
+Martini et al.
+(2009).
+The shortcoming of these formulas is that the momentum dependence is completely ignored.
+For this reason in Ref.
+Martini et al.
+(2009) Martini et al.
+introduced also a different approach that exploits the microscopic calculation of Alberico et al.
+Alberico et al.
+(1984) specifically aimed at the evaluation of the 2p2h contribution to the isospin spin-transverse response, measured in inclusive scattering.
+Their basic graphs are similar to those of Shimizu and Faessler Shimizu and Faessler (1980).
+In principle, this way of evaluation is definitely more satisfactory since the kinematical variables are correctly incorporated, but the results of Alberico et al.
+Alberico et al.
+(1984) are available only for a limited set of energy and momenta.
+Martini et al.
+have extended this range to cover the neutrino one through an approximate extrapolation.
+We refer to Ref.
+Martini et al.
+(2009) for the details.
+Martini et al.
+hence obtained, starting from the calculations of Alberico et al., “new” and responses entering in the new and components of the hadron tensor.
+For the part, which is not well covered in Ref.
+Alberico et al.
+(1984), Martini et al.
+have kept the previous parametrization, which already presents a proper dependence owing to the contribution of the in-medium width Oset and Salcedo (1987).
+To summarize, in the 2p2h sector, Martini et al.
+include NN correlations, MEC contributions, and NN correlations-MEC interference.
+Concerning the MEC, they decided to consider only the -MEC and to discard the other contributions from the explicit calculation because these other contributions are peculiar to the external probe and they want a “universal” spin-isospin 2p2h response, to use in (and deduced from) different processes, like electron scattering and pion absorption.
+The -MEC contribution is the dominant one in MEC, as shown, for example, in Fig.
+9 of De Pace et al.
+(2003).